# Zhaoxiang Zhang9, Yuanxing Zhang10, Ge Zhang3,7, Ke Xu1, Jie Fu11*, Wenhao Huang2,3

### v3

@@ Abstract
In this paper, we introduce MIO, a novel foundation model built on multimodal tokens, capable of understanding and generating speech, text, images, and videos in an end-to-end, autoregressive manner. While the emergence of large language models (LLMs) and multimodal large language models (MM-LLMs) propels advancements in artificial general intelligence through their versatile capabilities, they still lack true any-to-any understanding and generation. Recently, the release of GPT-4o has showcased the remarkable potential of any-to-any LLMs for complex real-world tasks, enabling omnidirectional input and output across images, speech, and text. However, it is closed-source and does not support the generation of multimodal interleaved sequences. To address this gap, we present MIO, which is trained on a mixture of discrete tokens across four modalities using causal multimodal modeling. MIO undergoes a four-stage training process: (1) alignment pre-training, (2) interleaved pre-training, (3) speech-enhanced pre-training, and (4) comprehensive supervised fine-tuning on diverse textual, visual, and speech tasks. Our experimental results indicate that MIO exhibits competitive, and in some cases superior, performance compared to previous dual-modal baselines, any-to-any model baselines, and even modality-specific baselines. Moreover, MIO demonstrates advanced capabilities inherent to its any-to-any feature, such as interleaved video-text generation, chain-of-visual-thought reasoning, visual guideline generation, instructional image editing, etc. Codes and models are available at https://github.com/MIO-Team/MIO.
@@ 1 Introduction
The advent of Large Language Models (LLMs) is commonly considered the dawn of artificial general intelligence (AGI) (OpenAI et al., 2023; Bubeck et al., 2023), given their generalist capabilities such as complex reasoning (Wei et al., 2022), role playing (Wang et al., 2023c), and creative writing (Wang et al., 2024a). However, original LLMs lack multimodal understanding capabilities. Consequently, numerous multimodal LLMs (MM-LLMs) have been proposed, allowing LLMs to understand images (Li et al., 2023b; Alayrac et al., 2022), audio (Borsos et al., 2023; Rubenstein et al., 2023; Tang et al., 2023; Das et al., 2024), and other modalities (Lyu et al., 2023; Zhang et al., 2023d; Moon et al., 2023). These MM-LLMs typically involve an external multimodal encoder, such as EVA-CLIP (Sun et al., 2023b) or CLAP (Elizalde et al., 2022), with an alignment module such as Q-Former (Li et al., 2023b) or MLP (Liu et al., 2023b) for multimodal understanding. These modules align non-textual-modality data features into the embedding space of the LLM backbone.
Table 1: The comparison between previous models and MIO (ours). I/O Consistency indicates whether the model ensures that the input and output representations for the same data remain consistent. Uni. Bi. SFT refers to whether the model undergoes a unified (Uni.) supervised fine-tuning (SFT) for both multimodal understanding and generation (Bi.=Bidirectional). Multi-Task SFT assesses whether the model undergoes a comprehensive SFT that includes diverse tasks, with at least visual question answering tasks. MM. Inter. Output evaluates whether the model supports the generation of multimodal interleaved (MM. Inter.) sequences. We refer readers to §1 for the definitions of the different modeling approaches.
Models Emu1
 (Sun et al., 2023c)
Emu2
@@
Modeling CICO CICO DIDO DIDO DIDO CIDO AR+Diff DIDO
@@
Another line of work involves building any-to-any and end-to-end MM-LLMs that can input and output non-textual modality data. Typically, there are four approaches: (1) Discrete-In-Discrete-Out (DIDO): Non-textual modality data is discretized using vector quantization techniques (van den Oord et al., 2017; Esser et al., 2020) and then fed into LLMs (Ge et al., 2023b; Zhan et al., 2024; Liu et al., 2024). (2) Continuous-In-Discrete-Out (CIDO): The LLM backbones intake densely encoded non-textual modality data features and generate their quantized representations (Diao et al., 2023; Team et al., 2023). (3) Continuous-In-Continuous-Out (CICO): The LLMs both understand and generate non-textual modality data in their densely encoded representations (Sun et al., 2023c; a; Dong et al., 2023; Zheng et al., 2023; Wu et al., 2023). (4) Autoregression + Diffusion (AR + Diff): The autoregressive and diffusion modeling are integrated in a unified LLM (Zhou et al., 2024; Xie et al., 2024; Li et al., 2024b). Although these works have succeeded in building MM-LLMs unifying understanding and generation, they exhibit some drawbacks, as illustrated in Table 1. For example, Emu1 (Sun et al., 2023c) and Emu2 (Sun et al., 2023a) explore the autoregressive modeling of three modalities: text, images, and videos. SEED-LLaMA (Ge et al., 2023b) proposes a new image quantizer aligned with LLMs’ embedding space and trains the MM-LLMs on images and videos. However, neither considers the speech modality, which is heterogeneous from visual modalities like videos and images. Although AnyGPT (Zhan et al., 2024) has explored settings involving four modalities, including text, image, speech, and music, it lacks video-related abilities, voice synthesis, and comprehensive multi-task supervised fine-tuning, leading to limited multimodal instruction-following and reasoning capabilities. Furthermore, AR + Diff approaches, such as Transfusion (Zhou et al., 2024), suffer from limited multimodal understanding capabilities because the multimodal inputs are noised for denoising modeling, and the image tokenizer used (i.e., VAE (Kingma & Welling, 2013)) is suitable for image generation rather than image understanding.
@@
Moreover, most of current MM-LLMs are typically dual-modal, combining text with another modality, such as images. Although previous works, such as Meta-Transformer (Zhang et al., 2023d) and Unified-IO 2 (Lu et al., 2023), have explored omni-multimodal understanding settings with more than two non-textual modalities, they still lag significantly behind their dual-modal counterparts, especially in terms of multimodal instruction-following capabilities. Moreover, these MM-LLMs are typically focused on understanding only, neglecting the important aspect of multimodal generation. Several works have enabled LLMs to call external tools to address this issue. For example, HuggingGPT (Shen et al., 2023) generates textual image descriptions for external diffusion models to synthesize images. GPT-4 (OpenAI et al., 2023) can utilize either an image generator like DALL-E 3 (Betker et al., 2024) or a text-to-speech (TTS) tool like Whisper (Radford et al., 2022) to support multimodal generation. [^1] https://openai.com/index/chatgpt-can-now-see-hear-and-speak/ However, these methods are not end-to-end, relying on the text modality as an interface.
@@
Recently, the release of GPT-4o has demonstrated the capabilities of any-to-any and end-to-end foundation models. [^2] https://openai.com/index/hello-gpt-4o/ It is the first foundational model to accept multimodal tokens as inputs and generate multimodal tokens within a unified model while also demonstrating strong abilities in complex multimodal instruction-following, reasoning, planning, and other generalist capabilities. Furthermore, as the continuous scaling up of LLMs in the community depletes high-quality language tokens, GPT-4o verifies a new source of data for LLM training: multimodal tokens. This approach suggests that the next generation AGI could derive more knowledge from multimodal tokens when language tokens are exhausted. However, GPT-4o is closed source and focuses primarily on end-to-end support for speech I/O, image I/O, 3D generation, and video understanding. Its recent open-source “alternatives”, such as VITA (Fu et al., 2024), still lack the ability to generate data of all supported modalities, particularly for the generation of multimodal interleaved sequences.
@@
To address the aforementioned issues, we introduce MIO (Multimodal Input and Output, or Multimodal Interleaved Output), the first open-source any-to-any foundation model that unifies multimodal understanding and generation across four modalities–text, image, speech (with voice), and video, while enabling the generation of multimodal interleaved sequences. Specifically, MIO is built on discrete multimodal tokens that capture both semantic representations through contrastive loss and low-level features via reconstruction loss (Ge et al., 2023a; Zhang et al., 2023b) from raw multimodal data. Due to the consistent data format shared with textual corpora, the model can treat non-textual modalities as “foreign languages”, allowing it to be trained with the next-token-prediction. Note that since the representation of an image remains the same whether it is used as an input or an output, our model flexibly supports multimodal interleaved sequence generation, where an image functions simultaneously for both understanding and generation. Moreover, we employ three-stage pre-training with an additional SFT stage to effectively train the model for modality scaling.
@@
Our experimental results show that MIO, trained on a mixture of four modalities, demonstrates competitive performance compared to its dual-modal counterparts and previous any-to-any multimodal language model baselines. Additionally, MIO is the first model to demonstrate interleaved video-text generation, chain-of-visual-thought reasoning, and other emergent abilities relying on any-to-any and multimodal interleaved output features (c.f.3.5).
@@ 2 Method
Firstly, we elaborate on our modeling approach, which supports multimodal token input and output, as well as causal language modeling (CausalLM), in §2.1. Secondly, we describe our three-stage pre-training procedures in §2.2. Thirdly, we provide details of our comprehensive supervised fine-tuning on diverse multimodal understanding and generation tasks in §2.3.
@@ 2.1 Modeling
In our work, we use SEED-Tokenizer (Ge et al., 2023a) as our image tokenizer and SpeechTokenizer (Zhang et al., 2023b) as our speech tokenizer. SEED-Tokenizer encodes images using a ViT (Dosovitskiy et al., 2021) derived from BLIP-2 (Li et al., 2023b), and then converts the encoded features into fewer tokens with causal semantics via Q-Former (Li et al., 2023b). These features are subsequently quantized into discrete tokens that are well-aligned with the language model backbone’s textual space. The codebook size for these discrete image tokens is 8192. SEED-Tokenizer transforms each image into a 224x224 resolution and quantizes it into 32 tokens. We use two special tokens, <image> and </image>, to indicate the start and end of the image tokens per image, respectively.
@@
In our preliminary experiments, we observed that text-to-speech generation (TTS) training is difficult to converge when using the alternating interleaving pattern because the noisy and random timbre tokens ($b_{1}c_{1}d_{1}$) tend to mislead the continuations. Moreover, the speech-to-text understanding (ASR) performance improves much more slowly during training with the alternating interleaving pattern due to the sparsity of semantic information in the timbre tokens. As a result, we drop the timbre tokens for speech understanding and use the sequential interleaving pattern for speech generation. We use <spch> and </spch> as special tokens to indicate the start and end of the speech token sequence.
@@
where $x_{t}$ represents the discrete multimodal tokens, and $\theta$ denotes the parameters of the LLM backbone. We use the pre-trained model, Yi-6B-Base (AI et al., 2024), for initialization.
Furthermore, to eliminate the computational inefficiency caused by <pad> tokens, we use the masked packing strategy (Lu et al., 2023; Liu et al., 2024; Dehghani et al., 2023). Specifically, the samples are concatenated along the sequence length dimension until the context window is full. Then, we construct the causal attention mask for the tokens of each sample and mask out all the tokens of the other samples.
@@ 2.2 Pre-Training
Pre-training Stage Stage I Stage II Stage III
Language-Only RefinedWeb RefinedWeb RefinedWeb
Image-Text Inter - OBELICS, MMC4-core-ff MMC4-core-ff
Video-Text Pair - WebVid-10M WebVid-10M
Speech-Text Pair Libriheavy Libriheavy Libriheavy
GPUs 128 A800-80GB 128 A800-80GB 8 A800-80GB
Training Steps 24,800 12,800 32,200
Batch Size 12:2:0:2 2:2:6:6 2:1:1:12
@@
In this stage, we extend the data used for pre-training to include image-text interleaved data (including video-text data) as a novel image-text dependency pattern. The image-text interleaving pattern has a different nature compared to pairing patterns. Although Li et al. (2023b) and Sun et al. (2023c) argued that interleaved image-text data mainly serves for multimodal in-context learning, we argue that it is also essential for context-aware image generation where images are generated based on specific context, rather than a precise description of the image content. For example, in image-text interleaved data, the text might serve as the image’s preceding or continuing context, rather than its description. This pattern significantly differs from the previous descriptive image generation demonstrated in image-text paired data, where images are generated based on precise and detailed text that clearly describe the content of the images (Team et al., 2023). Therefore, context-aware image generation is essential for tasks such as chain-of-visual-thought reasoning or visual storytelling (Team et al., 2023; Huang et al., 2016), where images are generated without textual descriptions. Due to the lack of benchmarks and evaluation metrics for context-aware image generation, we provide some demonstrations in §3.5 to showcase the potential of our model in visual storytelling, interleaved video-text generation, instructional image editing, chain-of-visual-thought reasoning, multimodal in-context learning, etc.
@@
To address this issue, we implement a three-stage strategy that gradually increases the proportion of speech tokens. In Stage I, speech-text data accounts for 12.5% of the training tokens, which rises to 37.5% in Stage II, and finally reaches 75.0% in Stage III. This incremental increase in the proportion of speech tokens ensures that the model’s performance in non-speech modalities is not compromised by the speech modality, while also allowing for the optimization of the model’s speech capabilities.
@@
We refer the reader to Appendix B for more details about the hyperparameters and prompt templates.

### v4

@@ Abstract
In this paper, we introduce MIO, a novel foundation model built on multimodal tokens, capable of understanding and generating speech, text, images, and videos in an end-to-end, autoregressive manner. While the emergence of large language models (LLMs) and multimodal large language models (MM-LLMs) propels advancements in artificial general intelligence through their versatile capabilities, they still lack true any-to-any understanding and generation. Recently, the release of GPT-4o has showcased the remarkable potential of any-to-any LLMs for complex real-world tasks, enabling omnidirectional input and output across images, speech, and text. However, it is closed-source and does not support the generation of multimodal interleaved sequences. To address this gap, we present MIO, which is trained on a mixture of discrete tokens across four modalities using causal multimodal modeling. MIO undergoes a four-stage training process: (1) alignment pre-training, (2) interleaved pre-training, (3) speech-enhanced pre-training, and (4) comprehensive supervised fine-tuning on diverse textual, visual, and speech tasks. Our experimental results indicate that MIO exhibits competitive, and in some cases superior, performance compared to previous dual-modal baselines, any-to-any model baselines, and even modality-specific baselines. Moreover, MIO demonstrates advanced capabilities inherent to its any-to-any feature, such as interleaved video-text generation, chain-of-visual-thought reasoning, visual guideline generation, instructional image editing, etc.
[Uncaptioned image] MIO: A Foundation Model on Multimodal Tokens
Zekun Wang1,3,13, King Zhu2,3, Chunpu Xu4, Wangchunshu Zhou5, Jiaheng Liu3,12, Yibo Zhang1, Jiashuo Wang4, Ning Shi6, Siyu Li3, Yizhi Li3,8, Haoran Que1, Zhaoxiang Zhang9, Yuanxing Zhang10,13, Ge Zhang3,7, Ke Xu1, Jie Fu11*, Wenhao Huang2,3thanks: Corresponding Authors. 1Beihang University; 201.AI; 3 M-A-P 4The Hong Kong Polytechnic University; 5AIWaves; 6University of Alberta; 7University of Waterloo; 8University of Manchester; 9 Chinese Academy of Sciences 10Peking University; 11Shanghai AI Lab; 12Nanjing University; 13Kuaishou Technology; zenmoore@buaa.edu.cn
@@ 1 Introduction
The advent of Large Language Models (LLMs) is commonly considered the dawn of artificial general intelligence (AGI) (OpenAI et al., 2023; Bubeck et al., 2023), given their generalist capabilities such as complex reasoning (Wei et al., 2022), role playing (Wang et al., 2023b), and creative writing (Wang et al., 2024a). However, original LLMs lack multimodal understanding capabilities. Consequently, numerous multimodal LLMs (MM-LLMs) have been proposed, allowing LLMs to understand images (Li et al., 2023b; Alayrac et al., 2022), audio (Borsos et al., 2023; Rubenstein et al., 2023; Tang et al., 2023; Das et al., 2024), and other modalities (Lyu et al., 2023; Zhang et al., 2023d; Moon et al., 2023). These MM-LLMs typically involve an external multimodal encoder, such as EVA-CLIP (Sun et al., 2023b) or CLAP (Elizalde et al., 2022), with an alignment module such as Q-Former (Li et al., 2023b) or MLP (Liu et al., 2023) for multimodal understanding. These modules align non-textual-modality data features into the embedding space of the LLM backbone.
Models Emu2
@@
Modeling CICO DIDO DIDO DIDO CIDO AR+Diff DIDO
Table 1: The comparison between previous models and MIO (ours). I/O Consistency indicates whether the model ensures that the input and output representations for the same data remain consistent. Uni. Bi. SFT refers to whether the model undergoes a unified (Uni.) supervised fine-tuning (SFT) for both multimodal understanding and generation (Bi.=Bidirectional). Multi-Task SFT assesses whether the model undergoes a comprehensive SFT that includes diverse tasks, with at least visual question answering tasks. MM. Inter. Output evaluates whether the model supports the generation of multimodal interleaved ($M$. Inter.) sequences. We refer readers to §1 for the definitions of the different modeling approaches.
@@
Another line of work involves building any-to-any and end-to-end MM-LLMs that can input and output non-textual modality data. Typically, there are four approaches: (1) Discrete-In-Discrete-Out (DIDO): Non-textual modality data is discretized using vector quantization techniques (van den Oord et al., 2017; Esser et al., 2020) and then fed into LLMs (Ge et al., 2023b; Zhan et al., 2024; Liu et al., 2024). (2) Continuous-In-Discrete-Out (CIDO): The LLM backbones intake densely encoded non-textual modality data features and generate their quantized representations (Diao et al., 2023; Team et al., 2023). (3) Continuous-In-Continuous-Out (CICO): The LLMs both understand and generate non-textual modality data in their densely encoded representations (Sun et al., 2023c, a; Dong et al., 2023; Zheng et al., 2023; Wu et al., 2023). (4) Autoregression + Diffusion (AR + Diff): The autoregressive and diffusion modeling are integrated in a unified LLM (Zhou et al., 2024; Xie et al., 2024; Li et al., 2024b). Although these works have succeeded in building MM-LLMs unifying understanding and generation, they exhibit some drawbacks, as illustrated in Table 1. For example, Emu1 (Sun et al., 2023c) and Emu2 (Sun et al., 2023a) explore the autoregressive modeling of three modalities: text, images, and videos. SEED-LLaMA (Ge et al., 2023b) proposes a new image quantizer aligned with LLMs’ embedding space and trains the MM-LLMs on images and videos. However, neither considers the speech modality, which is heterogeneous from visual modalities like videos and images. Although AnyGPT (Zhan et al., 2024) has explored settings involving four modalities, including text, image, speech, and music, it lacks video-related abilities, voice synthesis, and comprehensive multi-task supervised fine-tuning, leading to limited multimodal instruction-following and reasoning capabilities. Furthermore, AR + Diff approaches, such as Transfusion (Zhou et al., 2024), suffer from limited multimodal understanding capabilities because the multimodal inputs are noised for denoising modeling, and the image tokenizer used (VAE (Kingma and Welling, 2013)) is suitable for image generation rather than understanding.
@@
Moreover, most of current MM-LLMs are dual-modal, combining text with another modality, such as images. Although previous works, such as Meta-Transformer (Zhang et al., 2023d) and Unified-IO 2 (Lu et al., 2023), have explored omni-multimodal understanding settings with more than two non-textual modalities, they lag significantly behind their dual-modal counterparts, especially in terms of multimodal instruction-following capabilities. Moreover, these MM-LLMs are typically focused on understanding only, neglecting the important aspect of multimodal generation. Several works have enabled LLMs to call external tools to address this issue. For example, HuggingGPT (Shen et al., 2023) generates textual image descriptions for external diffusion models to synthesize images. GPT-4 (OpenAI et al., 2023) can utilize either an image generator like DALL-E 3 (Betker et al., 2024) or a text-to-speech (TTS) tool like Whisper (Radford et al., 2022) to support multimodal generation. [^1] https://openai.com/index/chatgpt-can-now-see-hear-and-speak/ However, these methods are not end-to-end, relying on the text modality as an interface.
@@
Recently, the release of GPT-4o has demonstrated the capabilities of any-to-any and end-to-end foundation models. [^2] https://openai.com/index/hello-gpt-4o/ It is the first foundational model to accept multimodal tokens as inputs and generate multimodal tokens in a unified model while also demonstrating strong abilities in complex multimodal instruction-following, reasoning, planning, and other generalist capabilities. Furthermore, as the scaling up of LLMs in the community depletes high-quality language tokens, GPT-4o verifies a new source of data for LLM training: multimodal tokens. This approach suggests that the next generation AGI could derive more knowledge from multimodal tokens when language tokens are exhausted. However, GPT-4o is closed source and focuses on end-to-end support for speech I/O, image I/O, 3D generation, and video understanding. Its open-source “alternatives”, such as VITA (Fu et al., 2024), still lack the ability to generate data of all supported modalities, particularly for the generation of multimodal interleaved sequences.
@@
To address the aforementioned issues, we introduce MIO (Multimodal Input and Output, or Multimodal Interleaved Output), the first open-source any-to-any foundation model that unifies multimodal understanding and generation across four modalities—text, image, speech (with voice), and video, while enabling the generation of multimodal interleaved sequences. Specifically, MIO is built on discrete multimodal tokens that capture both semantic representations through contrastive loss and low-level features via reconstruction loss (Ge et al., 2023a; Zhang et al., 2023b) from raw multimodal data. Due to the consistent data format shared with textual corpora, the model can treat non-textual modalities as “foreign languages”, allowing it to be trained with next-token prediction. Note that since the representation of an image remains the same whether it is used as an input or an output, our model supports multimodal interleaved sequence generation, where an image functions for both understanding and generation. Moreover, we employ three-stage pre-training with an additional SFT stage to effectively train the model for modality scaling.
@@
Our experimental results show that MIO, trained on a mixture of four modalities, demonstrates competitive performance compared to its dual-modal counterparts and previous any-to-any multimodal language model baselines. Additionally, MIO is the first model to demonstrate interleaved video-text generation, chain-of-visual-thought reasoning, and other emergent abilities relying on any-to-any and multimodal interleaved output features (c.f.E.5).
@@ 2 Method
Firstly, we elaborate on our modeling approach, which supports multimodal token input and output, as well as causal language modeling (CausalLM), in §2.1. Secondly, we describe our three-stage pre-training in §2.2. Thirdly, we provide details of our supervised fine-tuning on diverse multimodal understanding and generation tasks in §2.3.
@@ 2.1 Modeling
In our work, we use SEED-Tokenizer (Ge et al., 2023a) as our image tokenizer and SpeechTokenizer (Zhang et al., 2023b) as our speech tokenizer. SEED-Tokenizer encodes images using a ViT (Dosovitskiy et al., 2021) derived from BLIP-2 (Li et al., 2023b), and then converts the encoded features into fewer tokens with causal semantics via Q-Former (Li et al., 2023b). These features are subsequently quantized into discrete tokens that are well-aligned with the language model backbone’s textual space. The codebook size for these discrete image tokens is 8192. SEED-Tokenizer transforms each image into a 224x224 resolution and quantizes it into 32 tokens. We use two special tokens, <image> and </image>, to indicate the start and end of an image.
@@
In our preliminary experiments, we observed that text-to-speech generation (TTS) training is difficult to converge when using the alternating interleaving pattern because the noisy and random timbre tokens ($b_{1}c_{1}d_{1}$) tend to mislead the continuations. Moreover, the speech-to-text understanding (ASR) performance improves much more slowly during training with the alternating interleaving pattern due to the sparsity of semantic information in the timbre tokens. Thus, we drop the timbre tokens for speech understanding and use the sequential interleaving pattern for speech generation. We use <spch> and </spch> as special tokens to indicate the start and end of the speech token sequence.
@@
where $x_{t}$ denotes the discrete multimodal tokens, and $\theta$ denotes the LLM backbone parameters. We use Yi-6B-Base (AI et al., 2024) for initialization.
Furthermore, to eliminate the computational inefficiency caused by <pad> tokens, we use the masked packing strategy (Lu et al., 2023; Liu et al., 2024; Dehghani et al., 2023). The samples are concatenated along the sequence length until the context window is full. Then, we construct the causal attention mask for the tokens of each sample and mask out all the tokens of the other samples.
@@ 2.2 Pre-Training
In this stage, we extend the data used for pre-training to include image-text interleaved data (including video-text data) as a novel image-text dependency pattern. The image-text interleaving pattern has a different nature compared to pairing patterns. Although Li et al. (2023b) and Sun et al. (2023c) argued that interleaved image-text data mainly serves for multimodal in-context learning, we argue that it is also essential for context-aware image generation where images are generated based on specific context, rather than a precise description of the image content. For example, in image-text interleaved data, the text serves as the image’s preceding or continuing context, rather than its description. This pattern significantly differs from the previous descriptive image generation demonstrated in image-text paired data, where images are generated based on precise and detailed text that clearly describe the content of the images (Team et al., 2023). Therefore, context-aware image generation is essential for tasks like chain-of-visual-thought reasoning or visual storytelling (Team et al., 2023; Huang et al., 2016), where images are generated without textual descriptions. Due to the lack of benchmarks and evaluation metrics for context-aware image generation, we provide some demonstrations in §E.5 to showcase the potential of our model in visual storytelling, interleaved video-text generation, instructional image editing, chain-of-visual-thought reasoning, multimodal in-context learning, etc.
@@
To address this issue, we implement a three-stage strategy that gradually increases the proportion of speech tokens. In Stage I, speech-text data accounts for 12.5% of the training tokens, which rises to 37.5% in Stage II, and finally reaches 75.0% in Stage III. This incremental increase in the proportion of speech tokens ensures that the model’s performance in non-speech modalities is not compromised by the speech modality, while also allowing for the optimization of the model’s speech abilities.
@@
We refer the reader to Appendix B for details about the hyperparameters and prompt templates.